# DeepAccident: A Motion and Accident Prediction Benchmark for V2X Autonomous Driving

### 12 Dec 2023 v4

@@ Abstract
Tianqi Wang 1, Sukmin Kim 1, Wenxuan Ji 1, Enze Xie2 *,
Chongjian Ge 1, Junsong Chen 3, Zhenguo Li 2 Ping Luo1 *
1The University of Hong Kong, 2Huawei Noah’s Ark Lab, 3Dalian University of Technology
@@ Introduction
Accident generation details. For the two accident vehicles that we choose to collide with each other, we first calculate the intersection point of their planned trajectories and then perturb their initial positions to arrive at a similar time by dividing the arriving distance by the randomly chosen maximum speeds. The two vehicles designed to follow the accident vehicles have the same trajectories as the accident vehicles. Note that, the accident vehicles will also react to other surrounding vehicles to slow down via rule-based controllers when necessary, such that altering their time of arrival at accident sites. As a result, these accident vehicles may collide at diverse relative positions or angles and could potentially collide with other vehicles to increase trajectory diversity. Despite not being specifically designed, other accident scenarios occurred during data collection, such as vehicles hitting crossing pedestrians and colliding with lane-changing vehicles. These are included in the proposed DeepAccident dataset, with plans to incorporate more scenarios in future versions. Normal scenarios without collisions are also included in DeepAccident to enhance motion diversity. Each scenario will terminate when a collision occurs, the ego vehicle completes its planned trajectory, or the scenario time exceeds 10 seconds.
@@
V2X fusion module. We choose the five-agent V2X setting, and compare the performance of utilizing average pooling baseline and different state-of-the-art V2X fusion modules as V2XFormer’s fusion module. The V2X fusion methods used include DiscoNet (Li et al. 2021), V2X-ViT (Xu et al. 2022b), and CoBEVT (Xu et al. 2022a). As shown in Table 3, CoBEVT achieves the best performance in all three tasks and we will use this V2X fusion module for the following experiments.
@@
Robustness on pose error and latency. We test the robustness of V2XFormer with different V2X configurations against pose noise (Gaussian noise with a mean from 0.1m to 0.5m and a standard deviation of 0.02m), and delay latency (0.1s to 0.5s). As shown in Fig. 7, incorporating more V2X agents provides stronger robustness against pose error and communication latency. Nevertheless, communicating with more agents results in larger data transmission, which can subsequently increase time latency and hinder performance. Therefore, a trade-off is necessary when considering V2X configurations.

### 17 Dec 2023 v5

@@ Abstract
Tianqi Wang1, Sukmin Kim1, Ji Wenxuan1, Enze Xie2 *,
Chongjian Ge1, Junsong Chen3, Zhenguo Li2, Ping Luo1 *
1The University of Hong Kong
2Huawei Noah’s Ark Lab
3Dalian University of Technology
@@ Introduction
Accident generation details. For the two accident vehicles that we choose to collide with each other, we first calculate the intersection point of their planned trajectories and then perturb their initial positions to arrive at a similar time by dividing the arriving distance by the randomly chosen maximum speeds. The two vehicles designed to follow the accident vehicles have the same trajectories as the accident vehicles. Note that accident vehicles will also react to the surrounding traffic using rule-based controllers when necessary, thus altering their arrival times at the accident site. As a result, these accident vehicles may collide at diverse relative positions or angles and could potentially collide with other vehicles to increase trajectory diversity. Despite not being specifically designed, other accident scenarios occurred during data collection, such as vehicles hitting crossing pedestrians and colliding with lane-changing vehicles. These are included in the proposed DeepAccident dataset, with plans to incorporate more scenarios in future versions. Normal scenarios without collisions are also included in DeepAccident to enhance motion diversity. Each scenario will terminate when a collision occurs, the ego vehicle completes its planned trajectory, or the scenario time exceeds 10 seconds.
@@
V2X fusion module. We choose the five-agent V2X setting, and compare the performance of utilizing average pooling baseline and various state-of-the-art V2X fusion modules as V2XFormer’s fusion module. The V2X fusion methods used include DiscoNet (Li et al. 2021), V2X-ViT (Xu et al. 2022b), and CoBEVT (Xu et al. 2022a). As shown in Table 3, CoBEVT performs the best in all three tasks and we will use this V2X fusion module for the following experiments.
@@
Robustness on pose error and latency. We test the robustness of V2XFormer with different V2X configurations against pose noise (Gaussian noise with a mean from 0.1m to 0.5m and a standard deviation of 0.02m), and delay latency (0.1s to 0.5s). As shown in Figure 7, incorporating more V2X agents provides stronger robustness against pose error and communication latency. Nevertheless, communicating with more agents results in larger data transmission, which can subsequently increase time latency and hinder performance. Therefore, a trade-off is necessary when considering V2X configurations.
@@ Conclusion
Acknowledgement. This paper is partially supported by the General Research Fund of Hong Kong No.17200622.